# RFWave: Multi-band Rectified Flow for Audio Waveform Reconstruction

### v2

@@ Abstract
Recent advancements in generative modeling have significantly enhanced the reconstruction of audio waveforms from various representations. While diffusion models are adept at this task, they are hindered by latency issues due to their operation at the individual sample point level and the need for numerous sampling steps. In this study, we introduce RFWave, a cutting-edge multi-band Rectified Flow approach designed to reconstruct high-fidelity audio waveforms from Mel-spectrograms or discrete tokens. RFWave uniquely generates complex spectrograms and operates at the frame level, processing all subbands simultaneously to boost efficiency. Leveraging Rectified Flow, which targets a flat transport trajectory, RFWave achieves reconstruction with just 10 sampling steps. Our empirical evaluations show that RFWave not only provides outstanding reconstruction quality but also offers vastly superior computational efficiency, enabling audio generation at speeds up to 97 times faster than real-time on a GPU. An online demonstration is available at: https://rfwave-demo.github.io/rfwave/.
@@ 1 Introduction
Audio waveform reconstruction significantly enhances the digital interactions by enabling realistic voice and sound generation for diverse applications. This technology transforms low-dimensional features, derived from raw audio data, into perceptible sounds, improving the audio experience on various platforms such as virtual assistants and entertainment systems. The main challenges in this field focus on improving the quality of reconstructed audio and speeding up audio generation.
Prior to the widespread adoption of neural networks, audio waveform reconstruction methods such as STRAIGHT [1] and WORLD [2], based on speech signal processing techniques, were introduced. These methods reconstructed audio from low-dimensional features, delivering clear speech but sacrificing many details, rendering the output noticeably synthetic. However, deep learning has greatly enhanced audio waveform reconstruction models, evolving alongside neural network-based generative models.
Autoregressive models were first applied to the task of audio waveform reconstruction. WaveNet [3] notably advanced audio quality beyond prior signal processing methods, although it was hindered by slow generation speeds. Following this, WaveRNN [4] improved generation speed using a lightweight RNN and dual softmax layer. Furthermore, LPCNet [5] leveraged speech signal processing domain knowledge to reduce model complexity, enhancing both generation speed and audio clarity.
Nevertheless, the inherent slowness of autoregressive models, which generate sample points sequentially, sparked a demand for parallel generation. The rise of Generative Adversarial Networks (GAN) [6] inspired new research directions in audio waveform reconstruction. MelGAN [7], for example, utilized Mel-spectrograms as inputs to generate all corresponding waveform sample points in parallel, achieving speeds hundreds of times faster than real-time on GPUs. Models like Parallel WaveGAN [8] and HiFi-GAN [9] further refined the model architecture and discriminators, enhancing the quality of generated audio. Due to their relatively good performance and rapid generation speeds, GAN-based models quickly became mainstream in the industry. Considering that speech signals consist of a large number of sample points per second, modeling speech waveforms typically incorporates upsampling layers. A notable advancement in this aspect is the replacement of certain upsampling layers with the Inverse Short-Time Fourier Transform (ISTFT) in models like iSTFTNET [10], APNet2 [11], and Vocos [12]. This substitution not only simplifies the network architecture but also enhances computational efficiency, resulting in faster processing speeds while maintaining the quality of the speech.
Despite their advancements, GAN-based waveform reconstruction models face challenges such as the need for intricate discriminator designs and issues like instability and mode collapse. In response to these challenges, diffusion models for waveform reconstruction have been explored. For example, Diffwave [13] outperforms both WaveNet and WaveGAN [14]. PriorGrad [15] improves both the quality of speech and the speed of inference through the use of a data-dependent prior distribution, while FreGrad [16] simplifies the model and reduces denoising time by using Discrete Wavelet Transform (DWT).
Recently, neural codecs have gained increasing attention. These codecs offer superior compression rates compared to traditional signal processing methods, showing significant potential for applications in communication and storage. Additionally, they are capable of converting audio into discrete tokens, facilitating the integration of audio data with other modalities during the modeling process, as demonstrated by VALL-E [17] and UniAudio [18]. Typically, these neural codecs employ GAN-based decoder for waveform reconstruction, as seen in SoundStream [19] and EnCodec [20]. However, the recent Multi-Band Diffusion (MBD) [21] utilizes a diffusion model, which significantly enhances the quality and detail of the reconstructed waveform.
Despite producing relatively high-quality audio, diffusion-based waveform reconstruction has not gained widespread industry adoption due to the substantial computational costs and noticeable latency associated with its requirement for multiple sampling steps. In this paper, we propose RFWave, using Rectified Flow [22] for waveform reconstruction. Its inference speed can reach up to 97 times real-time on an NVIDIA GeForce RTX 4090 GPU. To our knowledge, RFWave stands as the fastest diffusion-based audio waveform reconstruction model, and it delivers superior audio quality. To accomplish its superior generation performance and computational efficiency, our main contributions are as follows:
@@
    By integrating the Rectified Flow and 3 enhanced loss functions – energy-balanced loss, overlap loss, and Short-Time Fourier Transform (STFT) loss – our model can reconstruct high-quality waveforms with a drastically reduced number of sampling steps.
@@
    We utilize a multi-band strategy, coupled with the high-efficiency ConvNeXtV2 [23] backbone, to generate different subbands concurrently. This not only assures audio quality by circumventing cumulative errors, but also boosts the synthesis speed.
@@ 2 Background
Rectified Flow [22] presents an innovative ODE-based framework for generative modeling and domain transfer. It introduces a method to learn a transport mapping that connects two distributions, $\pi_{0}$ and $\pi_{1}$ on $\mathbb{R}^{d}$, based on empirical observations:
@@
where $X_{t}=\phi(X_{0},X_{1},t)$ represents a time-differentiable interpolation between $X_{0}$ and $X_{1}$, with $\frac{\mathrm{d}}{\mathrm{d}t}X_{t}=\partial_{t}\phi(X_{0},X_{1},t)$. The $\gamma$ represents any coupling of $(\pi_{0},\pi_{1})$. An illustrative instance of $\gamma$ is the independent coupling $\gamma=\pi_{0}\times\pi_{1}$, which allows for empirical sampling based on separately observed data from $\pi_{0}$ and $\pi_{1}$. The authors recommended a simple choice of
@@
where the simulation is executed with a step size of $\epsilon=1/n$ over $n$ steps.
The velocity field has the capacity to incorporate conditional information, which is particularly essential in applications such as text-to-image generation and waveform reconstruction from Mel-spectrograms. Consequently, in such contexts, $v(Z_{t},t)$ in (2) is modified to $v(Z_{t},t\mid\mathcal{C})$, where $\mathcal{C}$ represents the conditional information pertinent to the corresponding $X_{1}$.
@@
Waveform reconstruction from complex spectrograms can be effectively achieved using the ISTFT. Notably, Vocos [12] and APNet2 [11], utilizing GANs as their model framework, estimate magnitude and phase spectrograms from the input Mel-spectrograms, which can be transformed to complex spectrograms effortlessly. Both models operate at the frame level, enabling them to achieve significantly faster inference speeds compared to HiFi-GAN [9], which uses multiple upsampling layers and operates at the level of waveform sample points. Morever, these models preserve the quality of the synthesized waveform, demonstrating their superiority in both speed and fidelity without a trade-off. In this paper, we directly estimate complex spectrograms using Rectified Flow and focus on frame-level operations, aiming to enhance both the efficiency and quality of our waveform synthesis process. Notably, the distributions of real and imaginary parts of the complex spectrograms appear more homogeneous in comparison to the distributions magnitude and phase.
@@
Both Multi-band MelGAN [24] and Multi-band Diffusion [21] employ multi-band strategies, albeit for different purposes within their respective frameworks. Multi-band MelGAN, specifically, uses Pseudo-Quadrature Mirror Filters (PQMF) [25] to divide frequency bands. This division results in each subband’s waveform being a fraction of the original waveform’s length, based on the number of subbands. By reshaping these subbands into feature dimensions and utilizing a unified backbone for modeling, Multi-band MelGAN is able to operate on considerably shorter signals. This strategy significantly enhances the efficiency of the model, leading to accelerated training and inference processes. Multi-band Diffusion utilizes an array of band-pass filters to separate the frequency bands and models each subband with a distinct model. This approach ensures that errors in one band do not negatively impact the others. In our research, we simplify the process of frequency band division by directly choosing the appropriate dimensions from the complex spectrograms. Furthermore, we enhance efficiency by modeling all subbands together in parallel with a single, unified model. This strategy improves the processing speed and also helps in reducing error accumulation across different subbands.
@@ 3 Method
Our model utilizes a multi-band Rectified Flow to directly predict the complex spectrogram. It operates at the STFT frame level and incorporates a highly efficient ConvNeXtV2 [23] backbone. With only 10 steps of sampling, the model is capable of producing high-quality waveforms. In this section, we present the structure of the multi-band Rectified Flow model, which operates in either the time or frequency domain. We also describe its corresponding normalization techniques and three enhanced loss functions.
(a) RFWave
(b) ConvNeXtV2 backbone
[…]
@@ 3.1 Multi-band Rectified Flow
The model structure is depicted in Figure 1(a). All frequency bands, each distinguished by a unique subband index, share the same model. The subbands of a given sample are grouped together into a single batch for processing, which facilitates simultaneous training or inference. This significantly reduces inference latency. Moreover, independently modeling the subbands reduces error accumulation. As discussed in [21], conditioning higher bands on lower ones can lead to an error accumulation, which means inaccuracies in the lower bands can adversely affect the higher bands during inference.
For each subband, the corresponding noise is fed into the ConvNeXtV2 backbone to predict the velocity conditioned on time $t$, the subband index, conditional input (the Mel-spectrogram or the EnCodec [20] tokens), and an optional EnCodec bandwidth index. The detailed structure of the ConvNeXtV2 backbone is shown in Figure 1(b). We employ Fourier features as described in [26]. The noise, Fourier features, and conditional inputs are concatenated and then passed through a linear layer, forming the input that is fed into a series of ConvNeXtV2 blocks. The sinusoidal time embedding, along with the optional EnCodec bandwidth index embedding, are element-wise added to the input of each ConvNeXtV2 block. Furthermore, the subband index is incorporated via an adaptive layer normalization module, which utilizes learnable embeddings as described in [12, 27]. The other components are identical to those within the ConvNeXtV2 architecture, details can be found in [23].
Our model, utilizing frame-level features, demonstrates greater memory efficiency compared to diffusion vocoders like PriorGrad that operate on waveform sample points. While PriorGrad [15] can train on 6-second [^1] The duration of an audio clip is calculated as the batch size multiplied by the number of segment frames and hop length, divided by the sampling rate. audio clips at 44.1 kHz within 30 GB of GPU memory, our model is capable of handling 177-second clips with the same memory resources.
@@
Our model is designed to function at the STFT frame level and possesses the capability to operate in the time domain. Assuming that the noise and velocity reside in the time domain, the use of STFT and ISTFT, as illustrated in Figure 1(a), becomes necessary. The dimensions of noise and velocity adhere to $[1,T]$, where $T$ represents the waveform length in sample points [^2] For simplicity, the batch dimension is not included in the discussion.. After the STFT operation, we extract the subbands by equally dividing the full-band complex spectrogram. Each subband $i$ is then processed as an independent sample. Subsequently, these processed subbands are merged back together before the ISTFT operation. The feature processed by the backbone thus has dimensions of $[d_{s},F]$, where $d_{s}$ denotes the dimension of a subband’s complex spectrum and $F$ the number of frames. The real and imaginary parts of each subband are interleaved to form a $d_{s}$-dimensional feature.
A white Gaussian noise signal has uniform energy distribution across all frequency bands. However, the energy profiles of various waveform types vary markedly among different frequency bands. For instance, the energy in a speech waveform exhibits an exponential decay with increasing frequency, whereas a music waveform tends to maintain a more consistent energy distribution across frequencies. These disparities pose challenges for training diffusion models. Consequently, it becomes advantageous to equalize the energy across waveform frequency bands [21]. In the time-domain model, a bank of Pseudo-Quadrature Mirror Filters (PQMF) is employed to decompose the input waveform into subbands. Subsequently, these subbands are equalized and then recombined to form the equalized waveform. The performance of the PQMF bank exhibits a modest enhancement compared to the array of band-pass filters employed in [21]. It’s important to note that the PQMF is solely utilized for waveform equalization and holds no association with the division of the complex spectrogram into subbands. For waveform equalization, mean-variance normalization is employed, utilizing the exponential moving average of mean and variance of each waveform subband computed during training. This approach ensures that the transformation can be effectively inverted using the same statistics.
@@
Assuming that noise and velocity reside in the frequency domain, the use of STFT and ISTFT, as depicted in Figure 1(a), becomes unnecessary. The subbands of noise and velocity are then directly extracted from their full-band feature by selecting the corresponding dimensions, resulting in a shape of $[d_{s},F]$. In the frequency-domain model, the waveform is transformed to a complex spectrogram without equalization. Subsequent processing involves the dimension-wise mean-variance normalization of the complex spectrogram.

### v3

@@ Abstract
Recent advancements in generative modeling have significantly enhanced the reconstruction of audio waveforms from various representations. While diffusion models are adept at this task, they are hindered by latency issues due to their operation at the individual sample point level and the need for numerous sampling steps. In this study, we introduce RFWave, a cutting-edge multi-band Rectified Flow approach designed to reconstruct high-fidelity audio waveforms from Mel-spectrograms or discrete acoustic tokens. RFWave uniquely generates complex spectrograms and operates at the frame level, processing all subbands simultaneously to boost efficiency. Leveraging Rectified Flow, which targets a straight transport trajectory, RFWave achieves reconstruction with just 10 sampling steps. Our empirical evaluations show that RFWave not only provides outstanding reconstruction quality but also offers vastly superior computational efficiency, enabling audio generation at speeds up to 160 times faster than real-time on a GPU. An online demonstration is available at: https://rfwave-demo.github.io/rfwave/.
@@ 1 Introduction
Audio waveform reconstruction significantly enhances the digital interactions by enabling realistic voice and sound generation for diverse applications. This technology transforms low-dimensional features, derived from raw audio data, into perceptible sounds, improving the audio experience on various platforms such as virtual assistants and entertainment systems. Autoregressive models and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) have been applied to this task, greatly advancing audio quality beyond traditional signal processing methods (Kawahara et al., 1999; Morise et al., 2016). Autoregressive methods, while effective, are hindered by slow generation speeds due to their sequential prediction of sample points (Oord et al., 2016; Kalchbrenner et al., 2018; Valin & Skoglund, 2019). In contrast, GANs predict sample points in parallel, resulting in faster generation speeds and maintaining high-quality output (Kumar et al., 2019; Yamamoto et al., 2020; Kong et al., 2020a; Siuzdak, 2023; Du et al., 2023). Consequently, GAN-based methods deliver impressive performance and are extensively utilized in real-world audio generation applications.
Despite the advancements, GAN-based waveform reconstruction models face challenges such as the necessity for complex discriminator designs and issues like instability or mode collapse (Thanh-Tung et al., 2018). In response, diffusion models for waveform reconstruction have been explored, offering stability during training and the ability to reconstruct high-quality waveforms (Chen et al., 2020; Kong et al., 2020b; gil Lee et al., 2022; Nguyen et al., 2024; Huang et al., 2022; Koizumi et al., 2022). However, these models are at least an order of magnitude slower compared to GANs. The slow generation speed in these diffusion-based waveform reconstruction models is primarily due to two factors: (1) the requirement of numerous sampling steps to achieve high-quality samples, and (2) the operation at the waveform sample point level. The latter often involves multiple upsampling operations to transition from frame rate resolution to sample rate resolution, increasing the sequence length and consequently leading to higher GPU memory usage and computational demands.
In this paper, we propose RFWave, a diffusion-type waveform reconstruction method designed to match the speed of GAN-based methods while maintaining the training stability and high sample quality of diffusion models. To overcome the challenge of slow sampling, we employ Rectified Flow (Liu et al., 2023; Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023), which connects data and noise along a straight line, thereby enhancing sampling efficiency. To address the GPU memory and computational demands of sample point-level modeling, our model operates at the level of Short-Time Fourier Transform (STFT) frames, enabling more efficient processing and reducing GPU memory usage. RFWave with only 10 sampling steps can generate high-quality audio, achieving an inference speed of up to 160 times real-time on an NVIDIA GeForce RTX 4090 GPU. Additionally, the incorporation of three enhanced loss functions and an optimized sampling strategy further elevates the overall quality of the reconstructed waveforms. To our knowledge, RFWave stands as the fastest diffusion-based audio waveform reconstruction model, and it delivers superior audio quality. Furthermore, it can reconstruct waveforms from Mel-spectrograms and discrete acoustic tokens, enhancing its versatility and applicability in various audio generation tasks. Our main contributions are as follows:
@@
    By integrating the Rectified Flow and 3 enhanced loss functions – energy-balanced loss, overlap loss, and STFT loss – our model can reconstruct high-quality waveforms with a drastically reduced number of sampling steps.
@@
    We utilize a multi-band strategy, coupled with the high-efficiency ConvNeXtV2 (Woo et al., 2023) backbone, to generate different subbands concurrently. This not only assures audio quality by circumventing cumulative errors, but also boosts the synthesis speed.
@@
  4. 4.
    We propose a technique for selecting sampling time points based on the straightness of the Rectified Flow transport trajectories, which enhances sample quality for free.
@@ 2 Background
Rectified Flow (Liu et al., 2023) presents an innovative Ordinary Differential Equation (ODE)-based framework for generative modeling and domain transfer. It introduces a method to learn a mapping that connects two distributions, $\pi_{0}$ and $\pi_{1}$ on $\mathbb{R}^{d}$, based on empirical observations:
@@
where $X_{t}=\phi(X_{0},X_{1},t)$ represents a time-differentiable interpolation between $X_{0}$ and $X_{1}$, with $\frac{\mathrm{d}}{\mathrm{d}t}X_{t}=\partial_{t}\phi(X_{0},X_{1},t)$. The $\gamma$ represents any coupling of $(\pi_{0},\pi_{1})$. An illustrative instance of $\gamma$ is the independent coupling $\gamma=\pi_{0}\times\pi_{1}$, which allows for empirical sampling based on separately observed data from $\pi_{0}$ and $\pi_{1}$. Liu et al. (2023) recommended a simple choice of
@@
where the simulation is executed with a step interval of $\epsilon=1/n$ over $n$ steps.
The velocity field has the capacity to incorporate conditional information, which is particularly essential in applications such as text-to-image generation and waveform reconstruction from compressed acoustic representation. Consequently, in such contexts, $v(Z_{t},t)$ in (2) is modified to $v(Z_{t},t\mid\mathcal{C})$, where $\mathcal{C}$ represents the conditional information pertinent to the corresponding $X_{1}$.
Diffusion Models for Audio Waveform Reconstruction
Diffusion models (Song et al., 2020; Ho et al., 2020) have become the de-facto choice for high-quality generation in the realm of generative models. Diffwave (Kong et al., 2020b) and WaveGrad (Chen et al., 2020) were pioneering efforts to reconstruct waveforms using diffusion models, achieving performance comparable to autoregressive models and GANs. PriorGrad (gil Lee et al., 2022) enhances both speech quality and inference speed by employing a data-dependent prior distribution. Meanwhile, FreGrad (Nguyen et al., 2024) simplifies the model and reduces denoising time through the use of Discrete Wavelet Transform (DWT). Multi-Band Diffusion (MBD) (Roman et al., 2023) leverages a diffusion model to reconstruct waveforms from discrete EnCodec (Défossez et al., 2022) tokens. Despite these advancements, the current fastest generation speed of these methods is only about 10 to 20 times faster than real-time, which limits their application for real-time use, especially when combined with large-scale Transformer-based acoustic models (Wang et al., 2023; Du et al., 2024). In contrast, our method achieves speeds up to 160 times faster than real-time, significantly enhancing real-time applicability.
@@
Waveform reconstruction from complex spectrograms can be effectively achieved using the ISTFT. Notably, Vocos (Siuzdak, 2023) and APNet2 (Du et al., 2023), utilizing GANs as their model framework, estimate magnitude and phase spectrograms from the input Mel-spectrograms, which can be transformed to complex spectrograms effortlessly. Both models operate at the frame level, enabling them to achieve significantly faster inference speeds compared to HiFi-GAN (Kong et al., 2020a), which uses multiple upsampling layers and operates at the level of waveform sample points. Morever, these models preserve the quality of the synthesized waveform, demonstrating their superiority in both speed and fidelity without a trade-off. In this paper, we directly estimate complex spectrograms using Rectified Flow and focus on frame-level operations, aiming to enhance both the efficiency and quality of our waveform synthesis process. Notably, the distributions of real and imaginary parts of the complex spectrograms appear more homogeneous in comparison to the distributions magnitude and phase.
@@
Both Multi-band MelGAN (Yang et al., 2021) and Multi-band Diffusion (Roman et al., 2023) employ multi-band strategies, albeit for different purposes within their respective frameworks. Multi-band MelGAN, specifically, uses Pseudo-Quadrature Mirror Filters (PQMF) (Johnston, 1980) to divide frequency bands. This division results in each subband’s waveform being a fraction of the original waveform’s length, based on the number of subbands. By reshaping these subbands into feature dimensions and utilizing a unified backbone for modeling, Multi-band MelGAN is able to operate on considerably shorter signals. This strategy significantly enhances the efficiency of the model, leading to accelerated training and inference processes. Multi-band Diffusion utilizes an array of band-pass filters to separate the frequency bands and models each subband with a distinct model. This approach ensures that errors in one band do not negatively impact the others. In our research, we simplify the process of frequency band division by directly choosing the appropriate dimensions from the complex spectrograms. Furthermore, we enhance efficiency by modeling all subbands together in parallel with a single, unified model. This strategy improves the processing speed and also helps in reducing error accumulation across different subbands.
@@ 3 Method
Our model utilizes a multi-band Rectified Flow to directly predict the complex spectrogram. It operates at the STFT frame level and incorporates a highly efficient ConvNeXtV2 (Woo et al., 2023) backbone. With only 10 steps of sampling, the model is capable of producing high-quality waveforms. In this section, we present the structure of the multi-band Rectified Flow model, which can operate in either the time or frequency domain. Additionally, we describe the corresponding normalization techniques, three enhanced loss functions, and the strategy for selecting sampling time points.
Figure 1: The overall structure for RFWave. $i_{sb}$ is the subband index, C is the conditional input, which can be Encodec token or Mel-spectrogram, and $i_{bw}$ is the EnCodec bandwidth index. Modules enclosed in a dashed box, as well as dashed arrows, are considered optional.
@@ 3.1 Multi-band Rectified Flow
The model structure is depicted in Figure 1. All frequency bands, each distinguished by a unique subband index, share the same model. The subbands of a given sample are grouped together into a single batch for processing, which facilitates simultaneous training or inference. This significantly reduces inference latency. Moreover, independently modeling the subbands reduces error accumulation. As discussed in (Roman et al., 2023), conditioning higher bands on lower ones can lead to an error accumulation, which means inaccuracies in the lower bands can adversely affect the higher bands during inference.
For each subband, the corresponding noisy sample ($X_{t}$) is fed into the ConvNeXtV2 backbone to predict the velocity ($v_{t}$) conditioned on time $t$, the subband index ($i_{sb}$), conditional input ($C$, the Mel-spectrogram or the EnCodec (Défossez et al., 2022) tokens), and an optional EnCodec bandwidth index ($i_{bw}$). The detailed structure of the ConvNeXtV2 backbone is shown in Figure 1. We employ Fourier features as described in (Kingma et al., 2021). The noisy sample, Fourier features, and conditional inputs are concatenated and then passed through a linear layer, forming the input that is fed into a series of ConvNeXtV2 blocks. The sinusoidal time embedding, along with the optional EnCodec bandwidth index embedding, are element-wise added to the input of each ConvNeXtV2 block. Furthermore, the subband index is incorporated via an adaptive layer normalization module, which utilizes learnable embeddings as described in (Siuzdak, 2023; Xu et al., 2019). The other components are identical to those within the ConvNeXtV2 architecture, details can be found in (Woo et al., 2023).
Our model, utilizing frame-level features, demonstrates greater memory efficiency compared to diffusion vocoders like PriorGrad that operate on waveform sample points. While PriorGrad (gil Lee et al., 2022) can train on 6-second [^1] The duration of an audio clip is calculated as the batch size multiplied by the number of segment frames and hop length, divided by the sampling rate. audio clips at 44.1 kHz within 30 GB of GPU memory, our model is capable of handling 177-second clips with the same memory resources.
@@
Our model is designed to function at the STFT frame level and possesses the capability to operate in the time domain. Assuming that the noisy sample and velocity reside in the time domain, the use of STFT and ISTFT, as illustrated in Figure 1, becomes necessary. The dimensions of noise and velocity adhere to $[1,T]$, where $T$ represents the waveform length in sample points [^2] For simplicity, the batch dimension is not included in the discussion.. After the STFT operation, we extract the subbands by equally dividing the full-band complex spectrogram. Each subband $i$ is then processed by the backbone as an independent sample. Subsequently, these processed subbands are merged back together before the ISTFT operation. The feature processed by the backbone thus has dimensions of $[d_{s},F]$, where $d_{s}$ denotes the dimension of a subband’s complex spectrum and $F$ the number of frames. The real and imaginary parts of each subband are interleaved to form a $d_{s}$-dimensional feature.
White Gaussian noise has uniform energy across frequency bands, but waveform energy profiles vary significantly. For example, speech energy decays exponentially with frequency, while music maintains a consistent distribution. These differences challenge diffusion model training, making energy equalization across bands beneficial (Roman et al., 2023). In the time-domain model, a bank of Pseudo-Quadrature Mirror Filters (PQMF) is employed to decompose the input waveform into subbands. Subsequently, these subbands are equalized and then recombined to form the equalized waveform. The performance of the PQMF bank exhibits a modest enhancement compared to the array of band-pass filters employed in (Roman et al., 2023). It’s important to note that the PQMF is solely utilized for waveform equalization and holds no association with the division of the complex spectrogram into subbands. For waveform equalization, mean-variance normalization is employed, utilizing the exponential moving average of mean and variance of each waveform subband computed during training. This approach ensures that the transformation can be effectively inverted using the same statistics.
@@
Assuming that noisy sample and velocity reside in the frequency domain, the use of STFT and ISTFT, as depicted in Figure 1, becomes unnecessary. The subbands of noisy sample and velocity are then directly extracted from their full-band feature by selecting the corresponding dimensions, resulting in a shape of $[d_{s},F]$. In the frequency-domain model, the waveform is transformed to a complex spectrogram without equalization. Subsequent processing involves the dimension-wise mean-variance normalization of the complex spectrogram.